# A Synergistic Framework of Nonlinear Acoustic Computing and Reinforcement Learning for Real-World Human–Robot Interaction

### v1

@@ Abstract
This paper introduces a novel framework integrating nonlinear acoustic computing and reinforcement learning to enhance advanced human–robot interaction under complex noise and reverberation. Leveraging physically informed wave equations (e.g., Westervelt, KZK), the approach captures higher-order phenomena such as harmonic generation and shock formation. By embedding these models in a reinforcement learning–driven control loop, the system adaptively optimizes key parameters (e.g., absorption, beamforming) to mitigate multipath interference and non-stationary noise. Experimental evaluations—covering far-field localization, weak signal detection, and multilingual speech recognition—demonstrate that this hybrid strategy surpasses traditional linear methods and purely data-driven baselines, achieving superior noise suppression, minimal latency, and robust accuracy in demanding real-world scenarios. The proposed system demonstrates broad application prospects in AI hardware, robot, machine audition, artificial audition, and brain-machine interfaces.
@@ 1   Introduction
Against this backdrop, this paper proposes an innovative framework that integrates nonlinear acoustic computing with reinforcement learning, creating a tightly coupled system capable of operating reliably in harsh acoustic environments. Nonlinear acoustic theories, derived from fundamental wave equations such as the Westervelt and Khokhlov–Zabolotskaya–Kuznetsov (KZK) equations, capture higher-order phenomena (harmonic generation, shock formation, waveform distortion, etc.) beyond the reach of simple linear models. These effects become pronounced when acoustic pressures are high or environmental boundary conditions induce notable nonlinearities, including strong reflections or concentrated energy focusing. Reinforcement learning augments these physically informed models by facilitating real-time parameter tuning—such as automatically adjusting attenuation coefficients and source localization strategies—to mitigate factors like multipath interference and unpredictable noise. Together, these components form an adaptive control loop, continuously refining both the physics-based model parameters and the signal processing front-end for robust performance.
@@
Center: A real-time auditory signal processing module encompasses adaptive filtering, source separation, and echo cancellation. Operating in tandem with the updated nonlinear model parameters, this subsystem addresses multiple acoustic challenges—ranging from abrupt changes in environmental noise to multi-speaker overlaps—empowering advanced human–robot interaction in noisy or highly reverberant spaces.
Right: In parallel, we incorporate a BMI-driven auditory enhancement unit. Neural and cortical signals are decoded and mapped onto parameters for dynamic speech reconstruction, offering individualized adjustments in intelligibility and perceptual quality. This integration lays the groundwork for next-generation brain–machine interfaces that combine advanced signal processing with user-specific customization.
@@
The rest of this paper is organized as follows. In Section 2, we describe the theoretical underpinnings of nonlinear acoustic computing, focusing on the derivations and practical aspects of the Westervelt and KZK equations, and propose a new reinforcement-learning-driven architecture for real-time parameter updating. Section 3 details the core algorithms, including AI-based noise reduction and echo cancellation (Section 3.1), a full-scenario sensor-array source localization approach (Section 3.2), a multilingual speech interaction pipeline capable of handling 66 languages (Section 3.3), and a multimodal large-model system supporting integrated acoustic–vision tasks (Section 3.4). Section 4 describes our experimental setup and results, highlighting metrics such as speech recognition accuracy, PESQ/MOS-LQO for noise suppression, real-time factor (RTF), and emotion/event detection, and comparing these against relevant baselines. Section 5 surveys various application scenarios, including AI headphones and smart speakers (Section 5.1), adaptive AI hearing aids (Section 5.2), multifunctional microphones and robot audition (Section 5.3), brain–machine interface hearing evaluations (Section 5.4), and novel in-vehicle acoustic systems (Section 5.5). Finally, Section 6 concludes the paper, outlining promising future directions such as multimodal fusion, edge–cloud co-deployment strategies, privacy-preserving learning, and the continuing evolution of next-generation human–robot interaction platforms.
@@ 2.1   Nonlinear Acoustic Modeling
Starting from the compressible Navier–Stokes equations under the assumptions of irrotational and lossless flow, one can obtain various forms of nonlinear acoustic wave equations through perturbation expansions. Let the total acoustic pressure be
@@ 2.1.1 The Westervelt Equation
By applying the second-order perturbation expansion to the governing Navier–Stokes equations under the assumptions stated above, one first arrives at an intermediate form:
@@
From this intermediate result, by normalizing and redefining parameters (absorbing $\epsilon$ into the new pressure variable) and dropping higher-order dissipative terms, one obtains the canonical Westervelt Equation:
$\frac{\partial^{2}p}{\partial t^{2}}\;-\;c^{2}\nabla^{2}p=\alpha\Bigl{(}\frac{\partial p}{\partial t}\Bigr{)}^{2},$ (2)
@@ 2.1.2 Derivation of the KZK Equation
For quasi-planar, directional beams in the far field, one must account for diffraction and absorption in addition to the nonlinearity captured by Equation (2). By introducing the paraxial approximation ($\partial/\partial z\ll\partial/\partial r$) and retarded time $\tau=t-z/c$, one arrives at the Khokhlov–Zabolotskaya–Kuznetsov (KZK) equation:
$\frac{\partial^{2}p}{\partial z\,\partial\tau}\;-\;\frac{c}{2}\,\nabla_{\perp}^{2}p=\frac{\alpha}{2\,\rho_{0}\,c^{3}}\,\frac{\partial^{3}(p^{2})}{\partial\tau^{3}}\;+\;\frac{\delta}{2\,c^{3}}\,\frac{\partial^{3}p}{\partial\tau^{3}},$ (3)
where $\nabla_{\perp}^{2}$ represents the transverse Laplacian with respect to the beam axis, and $\delta$ models thermoviscous absorption. With suitable scaling, one obtains a simplified form:
$\frac{\partial^{2}p}{\partial z\,\partial\tau}=\frac{c}{2}\,\nabla_{\perp}^{2}p\;+\;\frac{\alpha}{2\,\rho_{0}\,c^{3}}\,\frac{\partial^{3}(p^{2})}{\partial\tau^{3}}.$ (4)
In cylindrical coordinates $(r,\tau)$, an equivalent commonly cited formulation of the KZK Equation emerges:
$\frac{\partial^{2}p}{\partial\tau^{2}}\;-\;c^{2}\,\nabla^{2}p=\alpha\Bigl{(}\frac{\partial p}{\partial\tau}\Bigr{)}^{2}\;+\;\frac{1}{r}\,\frac{\partial}{\partial r}\!\Bigl{(}r\,\frac{\partial p}{\partial r}\Bigr{)},$ (5)
which more explicitly shows how nonlinearity, diffraction (through radial derivatives), and absorption (if the $\delta$-term is retained) combine to shape beam patterns and harmonic content. Equation (5) (the KZK Equation) can be specialized for different boundary conditions and is extensively used in modeling focused ultrasound beams, acoustic holography, and, by extension, certain far-field robotic audition or smart speaker array scenarios [2].
@@ 2.2.1 Deep Learning Synergy
    Feature Extraction: Multi-scale convolutional filters capture time–frequency patterns that characterize nonlinear distortion and reverberant tails.
@@ 3.4   Multimodal Large Model Technology System
Azero Acoustic Model: On the acoustic front, the Yuan Acoustic Model leverages petabyte-scale training data to push the boundaries of traditional speech emotion recognition. Even amidst high noise levels, it achieves an 80% accuracy rate in identifying emotional cues and supports a comprehensive catalog of 456 acoustic events. One secret sauce behind its success is robust real-time adaptation: the model continually recalibrates based on incoming signals, preserving resilience across chaotic noise environments. Such qualities have made it a fixture in cutting-edge smart home ecosystems and next-generation robot interfaces, cementing its reputation as a “one-stop shop” for dynamic audio intelligence.
@@ 4   Experimental Results and Performance Analysis
In high-noise environments, traditional speech recognition and noise suppression algorithms often struggle to simultaneously ensure accuracy and real-time performance. By contrast, the framework proposed in this paper effectively addresses these challenges, yielding competitive or superior results in four key areas: speech recognition rate, noise suppression capability, algorithmic computational latency, and emotion analysis. Furthermore, a comprehensive investigation of its performance under a variety of metrics highlights the system’s versatility and robustness in realistic deployment scenarios.
@@ 4.1   Noise Suppression Capability
Figure 6: PESQ MOS-LQO Quality Evaluation for Babble, Car, Street Noise.
@@ 4.5   Event and Emotion Recognition Capability
We further tested the AzeroASR event and emotion recognition model on a separate dataset (see Table LABEL:tab:comparison_azeroasr_otherasr), which includes comparisons against alternative systems [23]. The results reveal significant accuracy advantages in detecting nuances of speaker mood—even in the presence of 10 dB–30 dB background noise.
@@ 4.6   Language Understanding Capability
Livebench Overall 63.07 Computer Science 66.83 43
Biology 82.15 Economics 73.93
Business 66.67 Engineering 48.28
Chemistry 50.80 Average 64.53

### v2

# A Synergistic Framework of Nonlinear Acoustic Computing and Reinforcement Learning for Real-World Human-Robot Interaction

@@ Abstract
This paper introduces a novel framework integrating nonlinear acoustic computing and reinforcement learning to enhance advanced human-robot interaction under complex noise and reverberation. Leveraging physically informed wave equations (e.g., Westervelt, KZK), the approach captures higher-order phenomena such as harmonic generation and shock formation. By embedding these models in a reinforcement learning-driven control loop, the system adaptively optimizes key parameters (e.g., absorption, beamforming) to mitigate multipath interference and non-stationary noise. Experimental evaluations—covering far-field localization, weak signal detection, and multilingual speech recognition—demonstrate that this hybrid strategy surpasses traditional linear methods and purely data-driven baselines, achieving superior noise suppression, minimal latency, and robust accuracy in demanding real-world scenarios. The proposed system demonstrates broad application prospects in AI hardware, robot, machine audition, artificial audition, and brain-machine interfaces.
@@ 1   Introduction
Against this backdrop, this paper proposes an innovative framework that integrates nonlinear acoustic computing with reinforcement learning, creating a tightly coupled system capable of operating reliably in harsh acoustic environments. Nonlinear acoustic theories, derived from fundamental wave equations such as the Westervelt and Khokhlov-Zabolotskaya-Kuznetsov (KZK) equations, capture higher-order phenomena (harmonic generation, shock formation, waveform distortion, etc.) beyond the reach of simple linear models. These effects become pronounced when acoustic pressures are high or environmental boundary conditions induce notable nonlinearities, including strong reflections or concentrated energy focusing. Reinforcement learning augments these physically informed models by facilitating real-time parameter tuning—such as automatically adjusting attenuation coefficients and source localization strategies—to mitigate factors like multipath interference and unpredictable noise. Together, these components form an adaptive control loop, continuously refining both the physics-based model parameters and the signal processing front-end for robust performance.
@@
Center: A real-time auditory signal processing module encompasses adaptive filtering, source separation, and echo cancellation. Operating in tandem with the updated nonlinear model parameters, this subsystem addresses multiple acoustic challenges—ranging from abrupt changes in environmental noise to multi-speaker overlaps—empowering advanced human-robot interaction in noisy or highly reverberant spaces.
Right: In parallel, we incorporate a BMI-driven auditory enhancement unit. Neural and cortical signals are decoded and mapped onto parameters for dynamic speech reconstruction, offering individualized adjustments in intelligibility and perceptual quality. This integration lays the groundwork for next-generation brain-machine interfaces that combine advanced signal processing with user-specific customization.
@@
The rest of this paper is organized as follows. In Section 2, we describe the theoretical underpinnings of nonlinear acoustic computing, focusing on the derivations and practical aspects of the Westervelt and KZK equations, and propose a new reinforcement-learning-driven architecture for real-time parameter updating. Section 3 details the core algorithms, including AI-based noise reduction and echo cancellation (Section 3.1), a full-scenario sensor-array source localization approach (Section 3.2), a multilingual speech interaction pipeline capable of handling 66 languages (Section 3.3), and a multimodal large-model system supporting integrated acoustic-vision tasks (Section 3.4). Section 4 describes our experimental setup and results, highlighting metrics such as speech recognition accuracy, PESQ/MOS-LQO for noise suppression, real-time factor (RTF), and emotion/event detection, and comparing these against relevant baselines. Section 5 surveys various application scenarios, including AI headphones and smart speakers (Section 5.1), adaptive AI hearing aids (Section 5.2), multifunctional microphones and robot audition (Section 5.3), brain-machine interface hearing evaluations (Section 5.4), and novel in-vehicle acoustic systems (Section 5.5). Finally, Section 6 concludes the paper, outlining promising future directions such as multimodal fusion, real-world experience, and the continuing evolution of next-generation human-robot interaction platforms.
@@ 2.1   Nonlinear Acoustic Modeling
Starting from the compressible Navier-Stokes equations under the assumptions of irrotational and lossless flow, one can obtain various forms of nonlinear acoustic wave equations through perturbation expansions. Let the total acoustic pressure be
@@ 2.1.1 The Westervelt Equation
By applying the second-order perturbation expansion to the governing Navier-Stokes equations under the assumptions stated above, one first arrives at an intermediate form:
@@
From Equation (1), by normalizing and redefining parameters (absorbing $\epsilon$ into the new pressure variable) and dropping higher-order dissipative terms, one obtains the canonical Westervelt Equation:
$\frac{\partial^{2}p}{\partial t^{2}}\;-\;c^{2}\nabla^{2}p=\alpha\frac{\partial^{2}(p^{2})}{\partial t^{2}},$ (2)
@@ 2.1.2 Derivation of the KZK Equation
For quasi-planar, directional beams in the far field, one must account for nonlinearity, diffraction, and absorption. The Khokhlov-Zabolotskaya-Kuznetsov (KZK) equation is derived by applying the paraxial approximation ($\partial/\partial z\ll\partial/\partial r$) and using a retarded time frame $\tau=t-z/c$. This leads to the KZK equation:
$\frac{\partial^{2}p}{\partial z\,\partial\tau}=\frac{c}{2}\,\nabla_{\perp}^{2}p\;+\;\frac{\beta_{\text{NL}}}{2\,\rho_{0}\,c^{3}}\,\frac{\partial^{3}(p^{2})}{\partial\tau^{3}}\;+\;\frac{\delta}{2\,c^{3}}\,\frac{\partial^{3}p}{\partial\tau^{3}},$ (3)
where $\nabla_{\perp}^{2}$ represents the transverse Laplacian, $\beta_{\text{NL}}$ is the coefficient of nonlinearity (e.g., $1+B/2A$), $\rho_{0}$ is the ambient density, $c$ is the sound speed, and $\delta$ models thermoviscous absorption.
By neglecting absorption ($\delta=0$), a simplified, lossless form is obtained:
$\frac{\partial^{2}p}{\partial z\,\partial\tau}=\frac{c}{2}\,\nabla_{\perp}^{2}p\;+\;\frac{\beta_{\text{NL}}}{2\,\rho_{0}\,c^{3}}\,\frac{\partial^{3}(p^{2})}{\partial\tau^{3}}.$ (4)
For an axisymmetric beam in cylindrical coordinates $(r,z,\tau)$, where the transverse Laplacian is $\nabla_{\perp}^{2}p=\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial p}{\partial r}\right)$, the lossless KZK equation (from Eq. (4)) can be written explicitly as:
$\frac{\partial^{2}p}{\partial z\,\partial\tau}=\frac{c}{2r}\,\frac{\partial}{\partial r}\!\left(r\,\frac{\partial p}{\partial r}\right)\;+\;\frac{\beta_{\text{NL}}}{2\,\rho_{0}\,c^{3}}\,\frac{\partial^{3}(p^{2})}{\partial\tau^{3}}.$ (5)
This form explicitly shows how nonlinearity (via $\frac{\partial^{3}(p^{2})}{\partial\tau^{3}}$) and diffraction (via the radial derivative term) combine to shape beam patterns and harmonic content in axisymmetric systems. The KZK Equation (in its general forms (3), (4), or specialized form (5)) can be adapted for different boundary conditions and is extensively used in modeling focused acoustic holography, and related far-field scenarios [2].
@@ 2.2.1 Deep Learning Synergy
    Feature Extraction: Multi-scale convolutional filters capture time-frequency patterns that characterize nonlinear distortion and reverberant tails.
@@ 3.4   Multimodal Large Model Technology System
Azero Acoustic Model: On the acoustic front, the Yuan Acoustic Model leverages petabyte-scale training data to push the boundaries of traditional speech emotion recognition. Even amidst high noise levels, it achieves an 80% accuracy rate in identifying emotional cues and supports a comprehensive catalog of 450+ acoustic events. One secret sauce behind its success is robust real-time adaptation: the model continually recalibrates based on incoming signals, preserving resilience across chaotic noise environments. Such qualities have made it a fixture in cutting-edge smart home ecosystems and next-generation robot interfaces, cementing its reputation as a “one-stop shop” for dynamic audio intelligence.
@@ 4   Experimental Results and Performance Analysis
In high-noise environments, traditional speech recognition and noise suppression algorithms often struggle to simultaneously ensure accuracy and real-time performance. By contrast, the framework proposed in this paper effectively addresses these challenges, yielding competitive or superior results in four key areas: speech recognition rate, noise suppression capability, emotion analysis, and algorithmic computational latency. Furthermore, a comprehensive investigation of its performance under a variety of metrics highlights the system’s versatility and robustness in realistic deployment scenarios.
@@ 4.1   Noise Suppression Capability
Figure 6: PESQ MOS-LQO Quality Evaluation for Babble, Car, Street Noise
@@ 4.5   Event and Emotion Recognition Capability
We further tested the AzeroASR event and emotion recognition model on a separate dataset (see Table LABEL:tab:comparison_azeroasr_otherasr), which includes comparisons against alternative systems [20]. The results reveal significant accuracy advantages in detecting nuances of speaker mood—even in the presence of 10 dB-30 dB background noise.
@@ 4.6   Language Understanding Capability
Livebench Global Average 32.70 Reasoning Average 24.47 43
Coding Average 15.95 Mathematics Average 31.81
Data Analysis Average 33.95 Language Average 30.73
IF Average 59.31 All Item Average 32.70